Archive of Applied Mechanics manuscript No. (will be inserted by the editor)

First Author · Second Author

Insert your title here

Do you have a subtitle? If so, write it here

Received: date / Accepted: date

Abstract Insert your abstract here. Include up to five keywords.

Keywords First keyword · Second keyword · More

1 Introduction

Your text comes here. Separate text sections with

2 Section title

Citations [1] and [2] and [3].

2.1 Subsection title

as required. Don’t forget to give each section and subsection a unique label (see Sect. 2).

Paragraph headings Use paragraph headings as needed.

a^{2} + b^{2} = c^{2}  (1)

References

1. Author I (year) Article title. Journal Title-Abbreviated Vol: pp–pp

2. Author I, Smith J (year) Book title. Publisher, Place, pp numbers

3. Author I, Smith J (year) Paper title. In: Editor, A. (ed.) Proceedings Title, Location, Date, pages. Publisher, Place

F. Author first address Tel.: +123-45-678910 Fax: +123-45-678910 E-mail: fauthor@example.com

S. Author second address Fig. 1 Please write your figure caption here

Fig. 2 Please write your figure caption here Table 1 Please write your table caption here

first  second third
number number number
number number number